OpenStudy (clamin):

PLEASE HELP!! MEDAL!! Comparing factoring steps for different leading coefficients. Equation 1: 6x^2+10x -16=0 Equation 2: x^2 + 5/3 x - 8/3

OpenStudy (isaidavila):

What do you have to do?

OpenStudy (clamin):

i have to factor those equations and after that I have to factor it again with different leading coefficient @isaidavila

OpenStudy (anonymous):

\[6x^2+16x-6x-16=2x(3x+8)-2(3x+8)=(3x+8)(2x-2)=2(3x+8)(x-1)\]

OpenStudy (clamin):

i already factored it but my question is do i just have to choose any number for different leading coefficient?

OpenStudy (anonymous):

\[ 6x^2+10x -16=0 \iff x^2+5/3-8/3=0 \iff 6(x^2+5/3-8/3)=0\]

OpenStudy (anonymous):

You can multiply or divide both side by anything but 0. The same is true for the factored form.